# Fine-Grained Verifiers: Preference Modeling as Next-token Prediction in Vision-Language Alignment

### v3

@@ Abstract
The recent advancements in large language models (LLMs) and pre-trained vision models have accelerated the development of vision-language large models (VLLMs), enhancing the interaction between visual and linguistic modalities. Despite their notable success across various domains, VLLMs face challenges in modality alignment, which can lead to issues like hallucinations and unsafe content generation. Current alignment techniques often rely on coarse feedback and external datasets, limiting scalability and performance. In this paper, we propose FiSAO (Fine-Grained Self-Alignment Optimization), a novel self-alignment method that utilizes the model’s own visual encoder as a fine-grained verifier to improve vision-language alignment without the need for additional data. By leveraging token-level feedback from the vision encoder, FiSAO significantly improves vision-language alignment, even surpassing traditional preference tuning methods that require additional data. Through both theoretical analysis and experimental validation, we demonstrate that FiSAO effectively addresses the misalignment problem in VLLMs, marking the first instance of token-level rewards being applied to such models.
@@ 4 Experiment
Table 2: The performance of FiSAO  across all benchmarks is presented. The best result is bolded.
@@
Method MMEP MMEC SEED MMB MM-Vet SQAI POPE GQA Cap_val CHAIRS CHAIRI
LLaVA-1.5 1510.7 348.2 58.6 64.3 30.5 66.8 85.9 62.0 56.6 54.3 11.3
+ Vlfeedback 1432.7 321.8 59.3 64.0 31.2 66.2 83.7 63.2 54.8 40.3 13.2
+ Human-Prefer 1490.6 335.0 58.1 63.4 31.1 65.8 81.5 61.3 50.4 38.7 11.3
+ POVID 1452.8 325.3 60.2 64.9 31.8 68.8 86.9 61.7 57.3 35.2 8.3
+ FiSAO 1522.6 349.0 60.6 64.8 30.7 69.3 85.7 62.0 61.2 39.9 9.9
InstructBlip 1237.5 292.1 38.5 36.0 26.0 43.5 84.8 48.0 65.5 60.3 11.9
+ Vlfeedback 1241.3 298.9 40.4 37.7 26.6 44.6 78.5 47.7 64.0 56.5 9.7
+ Human-Prefer 1250.9 304.2 39.3 37.2 26.6 44.1 79.0 47.5 64.8 51.2 10.8
+ POVID 1255.1 301.8 38.3 37.2 26.3 43.4 84.6 48.3 66.5 51.5 10.5
+ FiSAO 1398.0 318.9 38.9 37.4 26.9 46.3 84.7 48.2 66.7 42.2 8.8
@@ 4.2 Experimental Results on Benckmarks (RQ1)
Comparison with Other Preference Tuning Approaches. As shown in Table 2, our method demonstrates clear advantages over other preference tuning approaches, which often require training reward models or incur high data costs. The superiority of FiSAO lies in its use of fine-grained verifier, which more effectively captures the intrinsic preferences of VLLMs and achieves stronger modality alignment between the pre-trained vison models and LLMs. Additionally, on the LLaVA backbone, FiSAO surpasses existing approaches, delivering an average performance improvement of 8.7%. This underscores FiSAO’s effectiveness in leveraging fine-grained token-level rewards to align visual and textual modalities seamlessly.
@@ 5 Related Work
Recent advancements in large language models (Brown et al., 2020; Touvron et al., 2023) and pre-trained vision models (Radford et al., 2021a) have enabled the creation of Vision-Large Language Models (VLLMs), which effectively integrate language and vision capabilities. These models have significantly improved automation in various fields, including medical applications (Liu et al., 2023b), autonomous driving (Zhou et al., 2023a), and embodied agents (Peng et al., 2023). The typical architecture of VLLMs involves aligning the embedding spaces of both modalities using techniques such as Qformer or fully connected layers (Zhu et al., 2023; Ye et al., 2023b; Li et al., 2023b). However, VLLMs face challenges in precise alignment due to independent pre-training of language and vision models, leading to safety concerns (Gong et al., 2023; Tu et al., 2023; Chen et al., ), hallucinations (Wang et al., 2023), and reasoning deficiencies (Ghosh et al., 2024). Traditional vision-language models (VLMs) have focused on image-text alignment through methods like co-attention frameworks (Lu et al., 2019), anchor points (Li et al., 2020), and contrastive learning (Radford et al., 2021b). Recently, alignment strategies can be classified into alignment from training data, which leverages high-quality datasets for supervised fine-tuning (SFT), and alignment from feedback, which involves fine-tuning based on human or AI feedback (Sun et al., 2023; Yu et al., 2023a; Zhou et al., 2024a; Li et al., 2023d; Zhao et al., 2023). Feedback-based methods often use Proximal Policy Optimization (PPO) (Sun et al., 2023) and Direct Preference Optimization (DPO) (Zhao et al., 2023; Li et al., 2023d; Chen et al., 2024a). Despite their potential, these methods face challenges such as high costs in dataset construction and need for external tools. Additionally, some approaches use sentence-level rewards, which do not fully leverage the text-generation capabilities that large language models (LLMs) are fundamentally designed for. By concentrating on assigning a numerical score to the entire instance, these methods overlook the VLLMs’ inherent capability to generate responses, including detailed reasoning steps. The detailed version is shown in Appendix A.4.

### v4

@@ Abstract
The recent advancements in large language models (LLMs) and pre-trained vision models have accelerated the development of vision-language large models (VLLMs), enhancing the interaction between visual and linguistic modalities. Despite their notable success across various domains, VLLMs face challenges in modality alignment, which can lead to issues like hallucinations and unsafe content generation. Current alignment techniques often rely on coarse feedback and external datasets, limiting scalability and performance. In this paper, we propose FiSAO (Fine-Grained Self-Alignment Optimization), a novel self-alignment method that utilizes the model’s own visual encoder as a fine-grained verifier to improve vision-language alignment without the need for additional data. By leveraging token-level feedback from the vision encoder, FiSAO significantly improves vision-language alignment, even surpassing traditional preference tuning methods that require additional data. Through both theoretical analysis and experimental validation, we demonstrate that FiSAO effectively addresses the misalignment problem in VLLMs, marking the first instance of token-level rewards being applied to such models. Our code is avaliable at https://github.com/gzcch/FISAO_ICLR.
@@ 4 Experiment
Table 2: The performance of FiSAO across all benchmarks. Bold indicates the best result and underline indicates the second-best result within each model group (LLaVA vs. InstructBlip). For CHAIRS and CHAIRI, smaller is better.
@@
Method MMEP MMEC SEED MM-Vet SQAI POPE GQA Cap_val CHAIRS CHAIRI
LLaVA-1.5 1510.7 348.2 58.6 30.5 66.8 85.9 62.0 56.6 54.3 11.3
+ Vlfeedback 1432.7 321.8 59.3 31.2 66.2 83.7 63.2 54.8 40.3 13.2
+ Human-Prefer 1490.6 335.0 58.1 31.1 65.8 81.5 61.3 50.4 38.7 11.3
+ POVID 1452.8 325.3 60.2 31.8 68.8 86.9 61.7 57.3 35.2 8.3
+ FiSAO 1522.6 349.0 60.6 30.5 69.3 85.7 62.0 61.2 39.9 9.9
InstructBlip 1237.5 292.1 38.5 26.0 43.5 84.8 48.0 65.5 60.3 11.9
+ Vlfeedback 1241.3 298.9 40.4 26.6 44.6 78.5 47.7 64.0 56.5 9.7
+ Human-Prefer 1250.9 304.2 39.3 26.6 44.1 79.0 47.5 64.8 51.2 10.8
+ POVID 1255.1 301.8 38.3 26.3 43.4 84.6 48.3 66.5 51.5 10.5
+ FiSAO 1305.3 308.9 40.0 26.9 45.4 83.7 48.2 66.7 42.2 8.8
@@ 4.2 Experimental Results on Benckmarks (RQ1)
Comparison with Other Preference Tuning Approaches. As shown in Table LABEL:tab:com, our method demonstrates clear advantages over other preference tuning approaches, which often require training reward models or incur high data costs. The superiority of FiSAO lies in its use of fine-grained verifier, which more effectively captures the intrinsic preferences of VLLMs and achieves stronger modality alignment between the pre-trained vison models and LLMs. Additionally, on the LLaVA backbone, FiSAO surpasses existing approaches, delivering an average performance improvement of 8.7%. This underscores FiSAO’s effectiveness in leveraging fine-grained token-level rewards to align visual and textual modalities seamlessly.
@@ 5 Related Work
Recent advancements in large language models (Brown et al., 2020; Liu et al., 2022; Touvron et al., 2023) and pre-trained vision models (Radford et al., 2021a) have enabled the creation of Vision-Large Language Models (VLLMs), which effectively integrate language and vision capabilities. These models have significantly improved automation in various fields, including medical applications (Liu et al., 2023b), recommendation (Sheng et al., 2024), autonomous driving (Zhou et al., 2023a), agent-based evaluation (Zheng et al., 2024),and embodied agents (Peng et al., 2023). The typical architecture of VLLMs involves aligning the embedding spaces of both modalities using techniques such as Qformer or fully connected layers (Zhu et al., 2023; Ye et al., 2023b; Li et al., 2023b). However, VLLMs face challenges in precise alignment due to independent pre-training of language and vision models, leading to safety concerns (Gong et al., 2023; Tu et al., 2023), hallucinations (Wang et al., 2023), and reasoning deficiencies (Ghosh et al., 2024). Traditional vision-language models (VLMs) have focused on image-text alignment through methods like co-attention frameworks (Lu et al., 2019), anchor points (Li et al., 2020), and contrastive learning (Radford et al., 2021b). Recently, alignment strategies can be classified into alignment from training data, which leverages high-quality datasets for supervised fine-tuning (SFT), and alignment from feedback, which involves fine-tuning based on human or AI feedback (Sun et al., 2023; Yu et al., 2023a; Zhou et al., 2024a; Li et al., 2023d; Zhao et al., 2023). Feedback-based methods often use Proximal Policy Optimization (PPO) (Sun et al., 2023) and Direct Preference Optimization (DPO) (Zhao et al., 2023; Li et al., 2023d; Chen et al., 2024a). Despite their potential, these methods face challenges such as high costs in dataset construction and need for external tools. Additionally, some approaches use sentence-level rewards, which do not fully leverage the text-generation capabilities that large language models (LLMs) are fundamentally designed for. By concentrating on assigning a numerical score to the entire instance, these methods overlook the VLLMs’ inherent capability to generate responses, including detailed reasoning steps. The detailed version is shown in Appendix A.4.
@@
Acknowledgement
This research is supported by the NExT Research Centre and Special Funding for Students’ Overseas Research Internships of University of Electronic Science and Technology of China (UESTC). We sincerely appreciate the reviewers and the AC for their valuable suggestions throughout the review process.
Ethics Statement
This paper aims to enhance vision-language alignment for Vision-Language Large Models (VLLMs) and do not obey the ICLR code of ethics.
Reproducibility Statement
All the results in this work are reproducible. We provide detailed settings for our experiments in Table 7.